OpenStudy (anonymous):

4sin^4 θ + 3sin^(2) θ-1=0

OpenStudy (anonymous):

let sin^2 theta=x so the problem reduces to, 4x^2+3x-1=0 so solving by middle term factor method, (4x-1)(x+1)=0 so, x= 1/4 as the other is not possible. so sin theta=sqrt(1/4)=(+,-)(1/2)

OpenStudy (anonymous):

do you understand the steps?